# On extension modules of finite homological dimension

### v1

@@ Abstract.
We explore the implications of the finiteness of homological dimensions for Ext modules, focusing on projective dimension, injective dimension, and their Gorenstein counterpart. In this direction, we establish several finiteness criteria for homological dimensions. Our results include the consequences of the finiteness of the Gorenstein (injective) dimension of the deficiency modules as well as a duality for certain Ext modules, all of finite (Gorenstein) projective dimension.
@@ 1. Introduction
After the astonishing ideas of Auslander, Buchsbaum, and Serre culminating in the characterization of the regular property of a Noetherian local ring by the finiteness of the projective dimension of its residual field, homological algebra has played a fundamental role in commutative ring theory. Indeed, this foundation enabled diverse characterizations of ring-theoretic properties via finiteness conditions on homological dimensions (e.g., projective, complete intersection, Gorenstein dimensions, and their injective versions), see for instance some classical references [1, 2, 6, 15, 29]. Establishing criteria for the finiteness of homological dimensions as well as their prescription – especially via Ext vanishing – is a very active topic in homological commutative algebra, containing a vast literature [1, 4, 6, 9, 12, 13, 18, 19, 21, 23, 27, 28, 26, 37, 36].
Recently, an approach to study the implications of vanishing of Ext is under the condition that certain Hom module has some finite homological dimension, as evidenced in [11, 12, 20, 28]. In addition, Ghosh-Puthenpurakal [18] and Zargar-Gheibi [37] studied the consequences of having Ext modules with finite injective dimension, deriving criteria for finiteness for the projective dimension of modules. Motivated by all, in this paper, we aim to study the consequences of having Ext modules with some finite homological dimension, focusing on the projective dimension, injective dimension, and their Gorenstein versions. Thus, the standard assumption in the present text does not revolve around the usual Ext vanishing hypothesis but rather the finiteness of its homological dimensions.
Zargar and Gheibi proved in [37] that for non-zero finitely generated modules $M$ and $N$ over a local ring $R$ such that $\operatorname{id}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $i$, the finiteness of the projective dimension of $M$ is equivalent to that of injective dimension of $N$. Then, what can one tell if one consider the Ext modules $\operatorname{Ext}_{R}^{i\geq 0}(M,N)$ with another finite homological dimension under the additional condition that $\operatorname{pd}_{R}(M)<\infty$ or $\operatorname{id}_{R}(N)<\infty$? Based on their approach employed in [37, Section 4], we obtain results concerning the finiteness of the homological dimension of Ext of complexes (see Theorem 3.3 and Theorem 3.6), obtaining consequently the following:
@@
Let $R$ be a local ring of dimension $d$, and let $M$ and $N$ be non-zero finitely generated $R$-modules. Let $\mathcal{H}$ be any of the homological dimensions $\operatorname{pd},\operatorname{G-dim},\operatorname{Gid}$, and suppose that $\mathcal{H}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq d-\operatorname{depth}(M)$.
@@
Theorem 1.1 not only gives criteria for detecting the finiteness of homological dimension of modules but also allows one to obtain Gorenstein criteria as verified in Corollary 3.5 and Corollary 3.8. In addition, it is worth noting that Theorem 1.1 says that the counterparts with Gorenstein dimensions of the results of Foxby in [14, Proposition 3.5(b)] are true. Furthermore, it is worth to note that item (2.a) allows to recover the celebrated result of Foxby given in [15, Corollary 4.4], and that item (2.c) gives a positive answer to the Gorenstein version of the question in [18, Question 4.5].
In our study, we also consider the deficiency modules, which were introduced by Schenzel [33]. They are an example of Ext modules, so what can one tell if all of the deficiency modules of a module have some finite homological dimension? Fiel and the first author [13] proved that the finiteness of the projective (resp. injective) dimension of all the deficiency modules of a finitely generated module $M$ over a local ring $R$ implies that $M$ has finite injective (resp. projective) dimension. We not only prove this result with a different approach, but also that its Gorenstein version is true.
@@
Let $R$ be a local ring with a dualizing complex and let $M$ be a finitely generated $R$-module.
@@
Over Cohen-Macaulay local rings admitting canonical module, we furnish a duality between certain Ext modules under the finiteness of some homological dimensions. Such a duality is also established for semidualizing modules (Theorem 5.5) and is applied in many directions throughout the paper.
@@
Assume $R$ Cohen-Macaulay local with canonical module $\omega$, and let $n$ be a non-negative integer. Let $\operatorname{H-dim}_{R}$ denote either the projective dimension or the Gorenstein dimension, and $\operatorname{Hid}_{R}$ the injective counterpart. Let $M$ and $N$ be finitely generated $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$. Then the following conditions are equivalent:
@@
This result leads to new finiteness criteria of homological dimension, for instance, we prove that if $R$ is a Cohen-Macaulay local ring of depth $t$ and, $M$ and $N$ are non-zero finitely generated $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then finiteness of the projective dimension $M$ is equivalent to the finiteness to that of $N$, see Corollary 5.9. Moreover, our theorem allows us to show that over Cohen-Macaulay local rings, the Ghosh-Puthenpurakal type formula [18, Theorem 3.5] is true in other classes of modules than G-perfect modules, and, consequently, furnishes a Gorensteinness criterion for such a ring, see Corollary 5.11.
The organization of this paper is as follows. In Section 2, since our study considers tools from the derived category, we give preliminaries about it. In Section 3, we establish results concerning Ext of complexes of finite homological dimension, prove Theorem 1.1, and derive Gorenstein criteria. In Section 4, we study the implications of having deficiency modules with some finite homological dimension. In Section 5, we focus our study on modules over Cohen-Macaulay local rings, prove Theorem 5.6, and derive several consequences. In Section 6, we discuss some questions motivated by results in Section 5.
Next, we establish our setup and provide simple examples of $\operatorname{Ext}$ modules with finite homological dimension to motivate our study.
@@
Throughout this paper, $R$ stands for a Noetherian (commutative unitary) ring. Whenever we assume that $R$ is local, its maximal ideal will be denoted by $\mathfrak{m}$ and its residue field by $k$; also, its depth is denoted by $t$. In that case, for a finitely generated $R$-module $M$, its $i$-th syzygy module and its transpose of Auslander are considered defined by a minimal free resolution of $M$, and denoted by $\Omega^{i}(M)$ and $\operatorname{Tr}_{R}(M)$, respectively. Also, for an $R$-module $M$, we let $\operatorname{H-dim}_{R}(M)$ stand for $\operatorname{pd}_{R}(M)$ or $\operatorname{G-dim}_{R}(M)$, and $\operatorname{Hid}_{R}(M)$ for the corresponding injective version. For any unexplained terminology, we refer to [5] and [10].
Example 1.4.
Assume $R$ is a local ring and let $\underline{x}=x_{1},\ldots,x_{n}$ be a regular sequence. Let $N$ be a finitely generated $R$-module such that $\underline{x}N=0$. Set $M=R/(\underline{x})$. Since the minimal free resolution of $M$ is the Koszul complex of the sequence $\underline{x}$, we have
$\operatorname{Ext}^{i}_{R}(L,X)\cong\operatorname{Ext}^{i}_{R}(R/(x_{1},\ldots,x_{i}),X)\cong X/(x_{1},\ldots,x_{i})X$
Thus $\operatorname{H-dim}_{R}(X)<\infty$ if and only if $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(L,X))<\infty$ for all $0\leq i\leq n$, and in that case, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(L,R))=i$, by the Auslander-Buchsbaum and Auslander-Bridge formulas.
Let $M$ be a finitely generated $R$-module. The grade of a finitely generated $R$-module $M$ is defined as $\operatorname{grade}_{R}(M)=\inf\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$. The module $M$ is said to be $\operatorname{H-}$perfect if $\operatorname{H-dim}_{R}(M)<\infty$ and $\operatorname{grade}(M)=\operatorname{H-dim}_{R}(M)$.
Example 1.5.
Let $M_{1},\ldots,M_{n}$ be H-perfect $R$-modules of grade $g_{1},\ldots,g_{n}$, respectively, and set $M=\bigoplus_{i=1}^{n}M_{i}$. Then
$\operatorname{Ext}^{g}_{R}(M,R)\cong\begin{cases}\operatorname{Ext}^{g_{i}}_{R}(M_{i},R),&\quad g=g_{i}\\ 0,&\quad\mbox{else}\end{cases}$
and $\operatorname{H-dim}_{R}(\operatorname{Ext}^{g_{i}}_{R}(M,R))=g_{i}$ by [3, Theorem 6.3].
An $R$-module $M$ is $p$-spherical if $\operatorname{pd}_{R}(M)\leq p$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $11\leq i\leq p-$ 1.
Example 1.6.
Based on the proof of [25, Proposition 6] we can construct spherical non-perfect modules as follows. We repeat the proof for convenience. Let $M$ be a finitely generated $R$-module with positive grade $n$, and let $F_{\bullet}$ be a free resolution of $M$. Consider the truncation
[…]
is a free resolution of $N$ and
[…]
is exact. This means that $N$ is $n$-spherical as well as a $(n+1)$th syzygy of $M$ (in the sense of [26]), and also $\operatorname{Ext}^{n}_{R}(N,R)\cong M$. It follows that whenever $M$ is such that $\operatorname{pd}_{R}(M)>\operatorname{grade}(M)=n>0$, there exists a $n$-spherical non-perfect $R$-module $N$ such that $\operatorname{Ext}^{n}_{R}(N,R)\cong M$. In particular, if $\operatorname{H-dim}_{R}(M)<\infty$ (resp. $\operatorname{Hid}_{R}(M)<\infty)$ then $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{n}(N,R)<\infty$).
It is noteworthy that the previous procedure works likewise if one considers resolutions by totally reflexive modules instead. In this case, $N$ would still be a non-G-perfect module which is a $(n+1)$th syzygy of $M$ (again in the sense of [26]) such that $\operatorname{G-dim}_{R}(N)\leq n$ and $\operatorname{Ext}^{i}_{R}(N,R)=0$ for all $i\neq 0,n$.
Example 1.7.
[…]
so that $\operatorname{Ext}^{2i+1}_{R}(M,M)\cong M$ for all $i$, which means that $\operatorname{CI-dim}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=0$ and $\operatorname{pd}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=\infty$.
2. Preliminaries on Derived Categories
In this section, we give some preliminaries concerning the derived category. For unexplained terminology, we refer to [6, Appendix], and for background on derived category theory, we additionally refer to [10].
2.1. Complexes and morphisms
All complexes considered in this section are descending. Let
$\boldsymbol{X}:\cdots\longrightarrow X_{i+1}\stackrel{{\scriptstyle\partial_{i+1}^{\boldsymbol{X}}}}{{\longrightarrow}}X_{i}\stackrel{{\scriptstyle\partial_{i}^{\boldsymbol{X}}}}{{\longrightarrow}}X_{i-1}\longrightarrow\cdots$
be a complex of $R$-modules.
If $X_{i}=0$ for all $i\gg 0$ (resp. for all $i\ll 0$), then $\boldsymbol{X}$ is said to be bounded to the left (resp. bounded to the right). If $\boldsymbol{X}$ is a bounded to the right and to the left, it is said to be bounded. We say that $\boldsymbol{X}$ is homologically bounded to the left (resp. homologically bounded to the right) if $\operatorname{H}_{i}(\boldsymbol{X})=0$ for all $i\gg 0$. If $\boldsymbol{X}$ is homologically bounded to the left and to the right, then $\boldsymbol{X}$ is said to be homologically bounded.
We recall that an morphism $f:\boldsymbol{X}\rightarrow\boldsymbol{Y}$ of complexes is a family $\{f_{i}:X_{i}\to Y_{i}\}_{i\in\mathbb{Z}}$ of homomorphism of $R$-modules such that $\partial_{i}^{\boldsymbol{Y}}\circ f_{i}=f_{i-1}\circ\partial_{i}^{\boldsymbol{X}}$. It is natural to define the composition of morphisms of complexes. The category of complexes over $R$ and homomorphism of $R$-modules is denoted by $\mathcal{C}(R)$.
Note that a morphism $f:\boldsymbol{X}\rightarrow\boldsymbol{Y}$ of complexes induces naturally homomorphisms $\operatorname{H}_{i}(f):\operatorname{H}_{i}(\boldsymbol{X})\rightarrow\operatorname{H}_{i}(\boldsymbol{Y})$. We say that $f$ is said to be a quasi-isomorphism if all the induced maps $H_{i}(f)$ are isomorphisms.
[…]
An $R$-module $M$ is identified naturally as a complex $\boldsymbol{X}$ such that $X_{0}=M$ and $X_{i}=0$ for all $i\neq 0$.
2.2. Notation of some full subcategories of $\mathcal{C}(R)$
We are considering the following full subcategories of $\mathcal{C}(R)$:
@@ 2.2. Notation of some full subcategories of $\mathcal{C}(R)$
    $\mathcal{C}_{\sqsubset}(R)$ (resp. $\mathcal{C}_{\sqsupset}(R)$) denotes the category of complexes bounded to the left (resp. right).
@@
    $\mathcal{C}_{(\sqsubset)}(R)$ (resp. $\mathcal{C}_{(\sqsupset)}(R)$) denotes the category of homologically bounded complexes to the left (resp. to the right).
  3. (3)
    $\mathcal{C}_{(\square)}(R)$ denotes the category of homologically bounded complexes.
  4. (4)
    $\mathcal{C}^{\rm P}(R)$ (resp. $\mathcal{C}^{\rm I}(R)$) denotes the category of complexes of $R$-modules of projective modules (resp. injective modules).
  5. (5)
    $\mathcal{C}^{(\rm f)}(R)$ denote the category of complexes with finitely generated homologies.
Moreover, for a subcategory $\mathcal{S}(R)\subseteq\mathcal{C}(R)$ we set $\mathcal{S}^{(\rm f)}(R)=\mathcal{S}(R)\cap\mathcal{C}^{(\rm f)}(R)$.
2.3. Shift
Let $n$ be an integer. For a complex $\boldsymbol{X},$ its $n$-fold shift $\Sigma^{n}\boldsymbol{X}$ is defined by:
$(\Sigma^{n}\boldsymbol{X})_{i}=X_{i-n}\,\,\text{ and }\,\,\partial_{i}^{\Sigma^{n}\boldsymbol{X}}=(-1)^{n}\partial_{i-n}^{\boldsymbol{X}}.$
2.4. Hom and tensor of complexes
Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes of $R$-modules. The Hom complex $\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})$ is defined by:
$\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})_{i}=\prod_{j\in\mathbb{Z}}\operatorname{Hom}_{R}(X_{j},Y_{i+j})$
and
$\partial_{i}^{\text{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})}((\psi_{j}))=(\partial_{j}^{\boldsymbol{Y}}\circ\psi_{j}-(-1)^{i}\psi_{p-1}\circ\partial_{j}^{\boldsymbol{X}})_{j}$
The tensor product complex $\boldsymbol{X}\otimes_{R}\boldsymbol{Y}$ is defined by:
$\left(\boldsymbol{X}\otimes_{R}\boldsymbol{Y}\right)_{i}=\bigoplus_{j\in\mathbb{Z}}X_{j}\otimes_{R}Y_{i-j}$
and
$\partial_{i}^{\boldsymbol{X}\otimes_{R}\boldsymbol{Y}}\left(m_{j}\otimes n_{i-j}\right)=\partial_{i}^{\boldsymbol{X}}\left(m_{j}\right)\otimes n_{i-j}+(-1)^{j}m_{j}\otimes\partial_{i-j}^{\boldsymbol{Y}}\left(n_{i-j}\right).$
2.5. Derived category
We denote by $\mathcal{D}(R)$ the derived category of $R$. This is the category of $R$-complexes localized at the class of all quasi-isomorphisms. Describing this category can be technical, so for more details, we refer the reader to [10]. Nevertheless, we mention the following facts:
  1. (1)
    The objects of $\mathcal{D}(R)$ are exactly the objects of $\mathcal{C}(R)$.
  2. (2)
    A morphism in $\mathcal{C}(R)$ is a morphism in $\mathcal{D}(R)$, and it is an isomorphism in $\mathcal{D}(R)$ if and only if it is a quasi-isomorphism in $\mathcal{C}(R)$.
By $\simeq$ we indicate isomorphism in $\mathcal{D}(R)$.
2.6. Derived functors
The symbols $\operatorname{\mathbf{R}Hom}(-,-)$ and $-\otimes_{R}^{\mathbf{L}}-$ denote, respectively, the right and left derived functors of the homomorphism and tensor product functors for complexes over $R$. They are bifunctors on the derived category of $R$, defined up to isomorphism in $\mathcal{D}(R)$.
The bifunctor $\operatorname{\mathbf{R}Hom}(-,-)$ satisfies the following: Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes.
  1. (1)
    If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{X}\simeq\boldsymbol{P}$, then $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{Hom}_{R}(\boldsymbol{P},\boldsymbol{Y})$.
  2. (2)
    If $\boldsymbol{Y}\in\mathcal{C}_{(\sqsubset)}(R)$ and if $\boldsymbol{I}$ is a complex in $\operatorname{C}_{\sqsubset}^{\rm I}(R)$ such that $\boldsymbol{I}\simeq\boldsymbol{Y}$, then $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{I})$.
Further, the bifunctor $-\otimes_{R}^{\mathbf{L}}-$ satisfies the following: Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes.
  1. (1)
    If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{X}\simeq\boldsymbol{P}$ , then $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}=\boldsymbol{P}\otimes_{R}\boldsymbol{Y}$.
  2. (2)
    If $\boldsymbol{Y}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{Y}\simeq\boldsymbol{P}$, then $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}=\boldsymbol{X}\otimes_{R}\boldsymbol{P}$.
If $\boldsymbol{X}$ and $\boldsymbol{Y}$ are complexes, we set
2.7. Standard isomorphisms
Let $\boldsymbol{X},\boldsymbol{Y}$ and $\boldsymbol{Z}$ be complexes. Then we have the following isomorphisms in $\mathcal{D}(R)$:
  1. (Com)
    $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}\simeq\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{L}$.
  2. (Asoc)
    $(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y})\otimes_{R}^{\mathbf{L}}\boldsymbol{Z}\simeq\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\left(\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{Z}\right).$
  3. (Adj)
    $\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y},\boldsymbol{Z}\right)\simeq\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{Y},\boldsymbol{Z})\right).$
  4. (Swap)
    $\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{Y},\boldsymbol{Z})\right)\simeq\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{Y},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Z})\right).$
2.8. Homological dimensions
(1) The projective dimension of a complex $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ is defined as:
$\operatorname{pd}_{R}(\boldsymbol{X})=\inf\{\sup\{i\in\mathbb{Z}:P_{i}\not=0\}:\boldsymbol{X}\simeq\boldsymbol{P}\in\mathcal{C}_{\sqsupset}^{\rm P}(R)\}.$
(2) The injective dimension of a complex $\boldsymbol{Y}\in\mathcal{C}_{(\sqsupset)}(R)$ is defined as:
$\operatorname{id}_{R}(\boldsymbol{Y})=\inf\{\sup\{i\in\mathbb{Z}:I_{-i}\not=0\}:\boldsymbol{Y}\simeq\boldsymbol{I}\in\mathcal{C}_{\sqsubset}^{\rm I}(R)\}.$
(3) A finitely generated $R$-module $G$ is totally reflexive if $G$ is reflexive and $\operatorname{Ext}_{R}^{i}(G,R)=\operatorname{Ext}_{R}^{i}(G^{\ast},R)=0$ for all $i>0$. By $\mathcal{C}^{\operatorname{G}}(R)$ we denote the category of complexes of totally reflexives $R$-modules. The Gorenstein dimension of a complex $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}^{({\rm f})}(R)$ is defined as
$\operatorname{G-dim}_{R}(\boldsymbol{X})=\inf\{\sup\{i\in\mathbb{Z}:G_{i}\not=0\}:\boldsymbol{X}\simeq\boldsymbol{G}\in\mathcal{C}_{\sqsupset}^{\rm G}(R)\}.$
(4) An $R$-module $M$ is said to be $G$-injective if there exists an exact complex of injective $R$-modules such that
$\boldsymbol{I}:\cdots\xrightarrow{d_{2}}{I_{2}}\xrightarrow{d_{1}}{I_{0}}\xrightarrow{d_{0}}I_{-1}\xrightarrow{d-1}\cdots$
such that $M\cong\ker(d_{0})$ and the complex $\operatorname{Hom}_{R}(J,\boldsymbol{I})$ is also exact for any injective $R$-module $J$. By $\mathcal{C}^{\operatorname{GI}}(R)$ we denote the category of complexes of Gorenstein injective modules. The Gorenstein injective dimension of a complex $\boldsymbol{Y}\in\mathcal{C}_{(\sqsubset)}(R)$ is defined as:
$\operatorname{Gid}_{R}(\boldsymbol{Y})=\inf\{\sup\{i\in\mathbb{Z}:I_{-i}\not=0\}:\boldsymbol{Y}\simeq\boldsymbol{I}\in\mathcal{C}_{\sqsubset}^{\rm GI}(R)\}.$

### v2

@@ Abstract.
We explore the implications of the finiteness of homological dimensions for Ext modules, focusing on projective dimension, injective dimension, and their Gorenstein counterpart. In this direction, we establish several finiteness criteria for homological dimensions. Under such finiteness conditions, our main result is a new duality for certain Ext modules when tensored by canonical modules.
@@ 1. Introduction
After the astonishing ideas of Auslander, Buchsbaum, and Serre culminating in the characterization of the regular property of a Noetherian local ring by the finiteness of the projective dimension of its residual field, homological algebra has played a fundamental role in commutative ring theory. Indeed, this foundation enabled diverse characterizations of ring-theoretic properties via finiteness conditions on homological dimensions (e.g., projective, complete intersection, Gorenstein dimensions, and their injective versions), see, for instance, some classical references [1, 2, 6, 14, 25]. Establishing criteria for the finiteness of homological dimensions as well as their prescription – especially via vanishing of Ext modules – is a very active topic in homological commutative algebra, containing a vast literature including but not limited to [1, 4, 6, 9, 11, 12, 15, 16, 18, 19, 23, 24, 22, 32, 31].
Recently, an approach to study the implications of vanishing of Ext is under the condition that certain Hom module has some finite homological dimension, as evidenced in [11, 17, 24]. In addition, Ghosh-Puthenpurakal [15] and Zargar-Gheibi [32] studied the consequences of having Ext modules with finite injective dimension, deriving criteria for finiteness for the projective dimension of modules. Motivated by all, in this paper, we aim to study the consequences of having Ext modules with some finite homological dimension, focusing on the projective dimension, injective dimension, and their Gorenstein versions. Thus, the standard assumption in the present text does not revolve around the usual Ext vanishing hypothesis but rather the finiteness of its homological dimensions.
Zargar and Gheibi proved in [32] that for non-zero finitely generated modules $M$ and $N$ over a Noetherian local ring $R$ such that $\operatorname{id}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $i$, the finiteness of the projective dimension of $M$ is equivalent to that of injective dimension of $N$. Then, what can one tell if one considers the Ext modules $\operatorname{Ext}_{R}^{i\geq 0}(M,N)$ with another finite homological dimension under the additional condition that $\operatorname{pd}_{R}(M)<\infty$ or $\operatorname{id}_{R}(N)<\infty$? We address this question in the following theorem.
@@
Let $R$ be a Noetherian local ring of dimension $d$, and let $M$ and $N$ be non-zero finitely generated $R$-modules. Let $\mathcal{H}$ be any of the homological dimensions $\operatorname{pd},\operatorname{G-dim},\operatorname{Gid}$, and suppose that $\mathcal{H}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq d-\operatorname{depth}(M)$.
@@
Theorem 1.1 not only gives criteria for detecting the finiteness of homological dimension of modules but also allows one to obtain Gorenstein criteria as verified in Corollary 2.4 and Corollary 2.7. In addition, it is worth noting that Theorem 1.1 says that the counterparts with Gorenstein dimensions of the results of Foxby in [13, Proposition 3.5(b)] are true. Furthermore, item (2.a) allows us to recover a celebrated result of Foxby given in [14, Corollary 4.4], and item (2.c) gives a positive answer to the Gorenstein version of the question in [15, Question 4.5].
In our study, we also consider deficiency modules, which were introduced by Schenzel [29]. Fiel and Holanda [12] proved that the finiteness of the projective (resp. injective) dimension of all deficiency modules of a finitely generated module $M$ over a local ring $R$ implies that $M$ has finite injective (resp. projective) dimension. We not only prove this result with a different approach, but also, we establish its Gorenstein version. The converse is assured for Cohen-Macaulay modules, extending Foxby [13].
@@
Let $R$ be a Noetherian local ring with a dualizing complex and let $M$ be a finitely generated $R$-module.
@@
Our main result is a new duality concerning the canonical module when tensored by certain Ext modules. Such a duality is also established for semidualizing modules (Theorem 4.5) and is applied in many directions throughout the paper.
@@
Assume $R$ is a Cohen-Macaulay local with canonical module $\omega$, and let $n$ be a non-negative integer. Let $\operatorname{H-dim}_{R}$ denote either $\operatorname{pd}$ or $\operatorname{G-dim}$, and $\operatorname{Hid}_{R}$ the injective counterpart. Let $M$ and $N$ be finitely generated $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$. Then the following conditions are equivalent:
@@
Using the behavior of the canonical module described in Theorem 1.3, we derive the following:
Corollary 1.4 (Corollary 4.8).
Let $R$ be a Cohen-Macaulay local ring of depth $t$ and, $M$ and $N$ be non-zero finitely generated $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{pd}_{R}(M)<\infty$ if and only if $\operatorname{pd}_{R}(N)<\infty$.
The organization of this paper is as follows. In Section 2, we establish results concerning Ext modules of complexes of finite homological dimension, and prove Theorem 1.1, and derive Gorenstein criteria. In Section 3, we study the implications of having deficiency modules with some finite homological dimension. In Section 4, we prove our main result, namely Theorem 1.3, and derive several consequences. In Section 5, we discuss some questions motivated by results in Section 4. Finally, in Section 6, we present some examples of Ext modules with finite homological dimension. Now, we establish our setup.
@@
Throughout this paper, $R$ stands for a Noetherian (commutative unitary) ring. Whenever we assume that $R$ is local, its maximal ideal will be denoted by $\mathfrak{m}$ and its residue field by $k$. The depth of $R$ is denoted by $t$. In that case, for a finitely generated $R$-module $M$, its $i$-th syzygy module and its transpose of Auslander are considered defined by a minimal free resolution of $M$, and denoted by $\Omega^{i}(M)$ and $\operatorname{Tr}_{R}(M)$, respectively.
We also consider the derived category of $R$. Unless, otherwise specified, we adopt the terminology used in [10]. In particular, we denote the projective and Gorenstein dimensions of an $R$-module $M$ respectively by $\operatorname{pd}_{R}(M)$ and $\operatorname{G-dim}_{R}(M)$, while the injective dimension and Gorenstein injective dimension of $M$ are respectively denoted by $\operatorname{id}_{R}(M)$ and $\operatorname{Gid}_{R}(M)$. In addition, for an $R$-module $M$, we let $\operatorname{H-dim}_{R}(M)$ stand for $\operatorname{pd}_{R}(M)$ or $\operatorname{G-dim}_{R}(M)$, and $\operatorname{Hid}_{R}(M)$ for the corresponding injective version.
If $\boldsymbol{X}$ and $\boldsymbol{Y}$ are complexes in $\mathcal{D}(R)$, we set
2. Finite homological dimensions of Ext of complexes
In this section, we derive results regarding the finiteness of homological dimension of Ext of complexes and conclude the corresponding results for Ext of modules.
Lemma 2.1.
Let $\boldsymbol{X}$ be a complex in $\mathcal{D}_{\square}(R).$
@@ 2. Finite homological dimensions of Ext of complexes
    If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i$, then $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$.
@@
    If $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i$, then $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Proof.
We prove item (1). This is done by induction on $s=\sup\boldsymbol{X}-\inf\boldsymbol{X}$. If $s=0$, then $\boldsymbol{X}\simeq\Sigma^{j}\operatorname{H}_{j}(\boldsymbol{X})$ for some $j$, and the conclusion follows.
Assume $s>0$ and let $u=\sup\boldsymbol{X}$. Note that there is a natural short exact sequence of complexes $0\to\boldsymbol{Y}\to\boldsymbol{X}\to\tau_{{}_{\subset u-1}}(\boldsymbol{X})\to 0,$ where $\tau_{{}_{\subset u-1}}(\boldsymbol{X})$ denotes the (t1)(t-1) soft left truncation of $\boldsymbol{X}$. In addition, note $\boldsymbol{X}\to\tau_{{}_{\subset u-1}}(\boldsymbol{X})$ induces an isomorphism in homologies in degree $\leq u-1$, and that $\operatorname{H}_{i}(\tau_{{}_{\subset u-1}}(\boldsymbol{X}))=0$ for i>u1i>u-1. Also, for $i\not=u$, we have that $\operatorname{H}_{i}(\boldsymbol{Y})=0$ and $\operatorname{H}_{u}(\boldsymbol{Y})\cong\operatorname{H}_{u}(\boldsymbol{X})$.
If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i$, note that $\boldsymbol{Y}$ and $\tau_{{}_{\subset t-1}(\boldsymbol{X})}$ satisfy the induction hypothesis, and therefore they have finite Gorenstein injective dimension. Thus, the conclusion follows from the fact that the class of complexes in $\mathcal{D}(R)$ with finite Gorenstein injective dimension has the two out three property [10, Proposition 9.2.15].
The proof of item (2) is similar using [10, Corollary 8.1.9 and Proposition 9.1.16]. ∎